OpenStudy (anonymous):

The measure of one of the acute angles in a right triangle is 63.7º. What is the measure of the other acute angle?

OpenStudy (anonymous):

180-90-63.7

OpenStudy (anonymous):

26.3 degrees

OpenStudy (apoorvk):

sum of all angles of a triangle is 180 degrees. let the unknown angle be x so, 90 + 63.7 + x = 180 or, x = 180 - 90 - 63.7 or, x = 90-63.7